A 4.0-Ω resistor, an 8.0-Ω resistor, and a 12-Ω resistor are connected in parallel with a 24-V battery. What are the equivalent resistance?
Answer #1

Consider R as the net resistance.
Since the connection is parallel,
1/R = 1/4 + 1/8 + 1/12
= (6 + 3 + 2)/24
= 11/24
Taking the reciprocal,
R = 24/11
= 2.18